OpenStudy (anonymous):

8 over x squared -4 plus 6 over 3x-6 solving quadratic equations :(?

OpenStudy (amistre64):

8 6 ------- + -------- like this? x^2 - 4 3x-6

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Does that equal something?

OpenStudy (anonymous):

but i dont know HOW to get the answer :/

OpenStudy (anonymous):

which is 2x +12 ------ x^2-4

OpenStudy (anonymous):

\[\frac{8}{x^2-4}+\frac{6}{3x-6}\] \[= \frac{8(3x-6) + 6(x^2-4)}{(3x-6)(x^2-4)}\] \[= \frac{8(3x-6) + 6(x+2)(x-2)}{(3x-6)(x^2-4)}\] \[= \frac{8(3x-6) + 2(x+2)3(x-2)}{(3x-6)(x^2-4)}\] \[= \frac{8(3x-6) + 2(x+2)(3x-6)}{(3x-6)(x^2-4)}\] \[= \frac{(3x-6)[8 + 2(x+2)]}{(3x-6)(x^2-4)}\] \[= \frac{[8 + 2(x+2)]}{x^2-4}\text{ , } x \ne 2\] \[= \frac{[8 + 2x+4)]}{x^2-4}\text{ , } x \ne 2\] \[= \frac{2x+12}{x^2-4}\text{ , } x \ne 2\]

OpenStudy (anonymous):

Actually x cannot be -2 either.